Consider three jobs – A, B, and C, with running times of 50
seconds, 30 seconds, and 70 seconds respectively. Assume context switching is
instantaneous.
a) Assume all jobs arrive at the same time, in the order A, then B, and then C. With respect
to FIFO, what is the average turnaround time and the average first-response time?
Answer #1

Ans

Average Response Time = 43.3 s

Average Turnaround Time = 93.3 s

In FIFO, processor is assigned according to their order.

Turnaround time = Time at which execution got completed - Arrival Time

Response Time is the time required to get the processor after arrival of process for the first time

= Time at processor assigned first time - Arrival time

Running Time of A = 50 seconds

After 50 seconds, execution of A completes

Turnaround Time of A = 50 s

At t=0, it gets the processor

Response Time of A = 0 s

After A, B starts executing

Running time of B = 30 second

After 30 seconds i.e at Time t = 50+30 = 80sec

Execution of B got completed.

Turnaround Time of B = 80 sec

B gets processor at t = 50 s

Response Time of B = 50 sec

Process C start executing at t = 80s

Execution Time = 70 s

After 70 s process gets completed, i.e at Time T = 80+70 = 150 s

It get completed at t = 150s

Turnaround Time = 150 s

It get processor at t = 80s

Response Time = 80 s

Average Turnaround Time = Sum of Turnaround Time/3

=(50+80+150)/3

= 280/3

= 93.3 s

Average Response Time = Sum of response Time / 3

= (0 + 50 + 80)/3

= 130/3

= 43.3s

If any doubt ask in the comments